# Leveraging Convolutional and Graph Networks for an Unsupervised Remote Sensing Labelling Tool

### v1

@@ 2.5 Graph Construction
Nodes for the graphs are determined by the Simple Linear Iterative Clustering (SLIC) algorithm (Achanta et al., 2012) where each segment is considered a new node. SLIC produces local homogenous parcels which for encoding geographical neighbours is ideal. We set the target number of segmentations for SLIC as $N=500$. The target number was chosen to over segment an image resulting in a roughly $1310m^{s}$ coverage for each segment, compromising computational cost in order to encode smaller features. This is only a target as SLIC can vary $N$ based on each sample as it splits or merges segments. The features of each node are $F_{i}=\{\bar{A}\}$. Where $\bar{A}$ is the mean values of each of the 64 activation maps from the CNN for each segment. Edges for each segment are determined by the $K$ geographically nearest segments. We tested on eight $K$ nearest neighbours. This work differs to previous work by Diao et al. (Diao et al., 2022) because we only include information from the CNN and rely on geographical neighbours to construct edges. The resulting representation is a Graph $G(V,E)$ with $N=V$ number of nodes with 64 features $F$ each. There are $K\times N$ unweighted edge connections between nodes, 8×500$\approx$ 8$\times$ 500≈ 8 × 500 edges.
@@ 2.6 Graph Neural Network
[…]
@@ 2.7 Graph Matching
We provide the user an interface that can switch between two modes. In the higher-level mode, users interact with $\times 256$ 256 pixel chips. In the lower-level mode, users interact with the segments. In both cases, the user interacts through a 2D UMAP projection of the manifold – Figure 2 (top) – and sees the relevant chips or segments in the view below – Figure 2 (bottom).
@@ 2.9.3 U-Net Evaluation
Part of our architecture is a U-Net that provides unsupervised features to feed into the GNN and graph matching. For this evaluation we compare our U-Net architecture to the state-of-the-art approaches of using ResNet-50 and GoogLeNet (Helber et al., 2019). The benchmarked approaches (Helber et al., 2019) were pre-trained on millions of images and then fine-tuned on EuroSat data. We utilise our pre-trained U-Net with two additional linear layers appended for classification. Our U-Net being pre-trained on only 141114111411 images, from a singular geographical region. If this test achieves an accuracy close to state-of-the-art we can attest that our U-Net, with some fine-tuning, has learnt rich features for the remaining parts of our pipeline. Successful application of the U-Net on a larger and more geographical diverse dataset also would demonstrate geo-generalisable weights.
@@ 3.1.1 Cluster exploration
In Figure 2 (top), the user interacts with the 2D manifold representation of the $\times 256$ 256 chip-level data by brushing selected points. The brushed points, highlighted in yellow, correspond to specific source images, which are simultaneously displayed in the image pane (Figure 2, bottom). This particular cluster reveals a patchwork of fields. The accompanying video illustrates the manifold’s spatial evolution by showing image chips transitioning through farmland, industrial zones, residential areas, increasing water bodies (e.g., lakes and sea), and other geographical features.
@@ 3.1.3 Segmentation Analysis (Figure 5)
Table 1: Respective classification accuracies for each model tested where C is the initial C means clustering number. Results labelled EuroSat where taken from (Helber et al., 2019).
@@ 3.4 Parameter Sensitivity
We used 8 edges during the graph generation phase. Here we compare results for $K\in\{4,8,12\}$. Our pipeline shows an improvement in for larger number of edges ($K$) in the neighbourhood (Table 4). Increasing the number of edges ($K$) increases the GCN training time and graph matching. For example, given the Hungarian algorithm’s complexity of $O(N^{3})$, where $N$ is the sum of all nodes in both graphs, the complexity for $K=8$ is $16^{3}$ whilst $K=12$ is $24^{3}$. Correspondingly, for the GCN training times for $K=8$ was 1.61.61.6 hours and $K=12$ was 6.66.66.6 hours. In our tests, larger $K$ produces more accuracy, but at significantly slower training times, and therefore we used $K=8$ as a good balance.
@@
Table 5: Comparison of varying SLIC parameter $N$, used during the graph construction phase Section 2.5.
Here we report values of $N\in{200,500,800}$ for the number of SLIC segments. Our initial choice, $N=500$, was to cover an approximate area of $1300m^{2}$ for ease of interpretability and efficiency when visualised for the user (the mean average size of segments is 256×256÷500=131256$\times$ 256$\div$ 500=131256 × 256 ÷ 500 = 131 — approximately 11 by 11 pixels and each pixel is $10m^{2}$ in Sentinel-2). A smaller area allows finer labelling at the cost of speed, namely projecting and labelling more points. This choice for $N$ was also optimal when considering local neighbourhood encoding using the different measures, see Table 5. For smaller $N$ values, we see that comparing segments via GLCM shows better results however, over all measures $N=500$ is the most accurate.
@@ Acknowledgements
This research was funded by EPSRC grant number EP/S021892/1 and UKHO. For the purpose of open access the authors have applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript version arising from this submission.

### v2

@@ 2.5 Graph Construction
Nodes for the graphs are determined by the Simple Linear Iterative Clustering (SLIC) algorithm (Achanta et al., 2012) where each segment is considered a new node. SLIC produces local homogenous parcels which for encoding geographical neighbours is ideal. We set the target number of segmentations for SLIC as $N=500$. This choice was made to over-segment the image, resulting in a roughly $1310\,m^{2}$ coverage for each segment, which provides finer granularity while balancing computational cost. Although SLIC dynamically adjusts $N$ based on each sample as it splits or merges segments, we selected $N=500$ as a practical compromise between segment size and efficiency. This decision was supported by sensitivity analysis (Section 3.4), where $N=500$ consistently achieved the best trade-off between accuracy and performance compared to $N=200$ and $N=800$. We used SLIC for super-pixel generation due to its computational efficiency and ability to produce spatially compact, homogeneous segments, which are well-suited for encoding geographical neighbourhoods in graph structures. SLIC is widely used in both computer vision and remote sensing because it offers fine control over segment size and compactness while maintaining low memory overhead, an important consideration for large Sentinel-2 tiles. Although Sentinel-2 imagery contains multiple spectral channels, we applied SLIC to the RGB bands only. This choice reflects a trade-off between segmentation quality and computational cost: RGB channels capture dominant spatial and textural patterns for defining segment boundaries, while the full multi-band spectral information is incorporated later during CNN feature extraction.
The features of each node are $F_{i}=\{\bar{A}\}$. Where $\bar{A}$ is the mean values of each of the 64 activation maps from the CNN for each segment. Edges for each segment are determined by connecting to the $K$ geographically nearest segments, where $K=8$. This strategy was chosen to preserve spatial context, which is critical for remote-sensing semantics: for example, distinguishing a water segment surrounded by ocean from one adjacent to urban areas. Using feature similarity based K-NN would risk linking distant regions with similar textures, losing neighbourhood information that our pipeline relies on for context-aware embeddings. Geographical adjacency also offers computational advantages, as edge connections remain fixed and do not require iterative recomputation during training. The choice of $K=8$ was supported by sensitivity analysis (Section 3.4), where this value achieved the best balance between accuracy and efficiency. This work differs from previous approaches by Diao et al. (Diao et al., 2022) because we only include information from the CNN and rely on geographical neighbours to construct edges. The resulting representation is a graph $G(V,E)$ with $N=V$ nodes, each with 64 features $F$. There are $K\times N$ unweighted edge connections between nodes, approximately $8\times 500$ edges.
@@ 2.6 Graph Neural Network
where $k$ is the number of attention heads, $\big\|$ denotes vector concatenation, $\sigma(\cdot)$ is an activation function, $N$ is the Neighbourhood of edges to $i$, $\mathbf{W}^{k}$ is a matrix of parameters for the $k$-th attention head and $\alpha_{ij}$ are attention coefficients defined by the following:
@@ 2.7 Graph Matching
We provide the user an interface that can switch between two modes. In the higher-level mode, users interact with $256\times 256$ pixel chips. In the lower-level mode, users interact with the segments. In both cases, the user interacts through a 2D UMAP projection of the manifold – Figure 2 (top) – and sees the relevant chips or segments in the view below – Figure 2 (bottom).
@@
Front End (User Interface).
The front end provides the interactive 2D manifold view (UMAP) at two levels: (i) chip-level points, and (ii) segment-level points for selected images. Users can pan, zoom, and brush rectangular selections. Selected points are reflected in a linked image panel that renders the corresponding chip thumbnails or super-pixel masks. For performance, the user interface is written using DirectX 11 and C++. The real-time interaction (in the accompanying video) was captured whilst using an NVIDIA RTX 3090.
Back End (Model and Data Handling).
All model training and inference are implemented in PyTorch with GPU acceleration via CUDA. The offline pipeline includes U-Net training for feature extraction, GNN training for context-aware embeddings, and UMAP projection for dimensionality reduction. These activation maps, graph embeddings, and 2D coordinates, are saved to files and accessed by the front-end interface for interactive exploration. During interaction, the tool reads these precomputed files to render chip thumbnails, segmentation masks, and manifold positions in real time. This design separates heavy batch computation (performed offline on an NVIDIA RTX 3090 GPU) from lightweight rendering and selection operations.
@@ 2.9.3 U-Net Evaluation
Part of our architecture is a U-Net that provides unsupervised features to feed into the GNN and graph matching. For this evaluation we compare our U-Net architecture to the state-of-the-art approaches of using ResNet-50 and GoogLeNet (Helber et al., 2019). The benchmarked approaches (Helber et al., 2019) were pre-trained on millions of images and then fine-tuned on EuroSat data. We utilise our pre-trained U-Net with two additional linear layers appended for classification. Our U-Net being pre-trained on only $1411$ images, from a singular geographical region. If this test achieves an accuracy close to state-of-the-art we can attest that our U-Net, with some fine-tuning, has learnt rich features for the remaining parts of our pipeline. Successful application of the U-Net on a larger and more geographical diverse dataset also would demonstrate geo-generalisable weights.
@@ 3.1.1 Cluster exploration
In Figure 2 (top), the user interacts with the 2D manifold representation of the $256\times 256$ chip-level data by brushing selected points. The brushed points, highlighted in yellow, correspond to specific source images, which are simultaneously displayed in the image pane (Figure 2, bottom). This particular cluster reveals a patchwork of fields. The accompanying video illustrates the manifold’s spatial evolution by showing image chips transitioning through farmland, industrial zones, residential areas, increasing water bodies (e.g., lakes and sea), and other geographical features.
@@ 3.1.3 Segmentation Analysis (Figure 5)
Table 1: Classification accuracies for each model tested, where $C$ is the initial Fuzzy C-Means clustering number. EuroSat results are taken from (Helber et al., 2019). These results serve as a validation of the U-Net component rather than the primary contribution, confirming its ability to learn generalisable features for use in our unsupervised labelling pipeline.
@@
It is important to note that Table 1 is not intended as a primary result of this work but as a validation step to demonstrate that our U-Net, trained in an unsupervised manner, learns generalisable features suitable for integration into the full pipeline.
@@ 3.4 Parameter Sensitivity
We used 8 edges during the graph generation phase. Here we compare results for $K\in\{4,8,12\}$. Our pipeline shows an improvement in for larger number of edges ($K$) in the neighbourhood (Table 4). Increasing the number of edges ($K$) increases the GCN training time and graph matching. For example, given the Hungarian algorithm’s complexity of $O(N^{3})$, where $N$ is the sum of all nodes in both graphs, the complexity for $K=8$ is $16^{3}$ whilst $K=12$ is $24^{3}$. Correspondingly, for the GCN training times for $K=8$ was $1.6$ hours and $K=12$ was $6.6$ hours. In our tests, larger $K$ produces more accuracy, but at significantly slower training times, and therefore we used $K=8$ as a good balance.
Table 5: Comparison of varying SLIC parameter NN, used during the graph construction phase Section 2.5.
@@
Here we report values of $N\in{200,500,800}$ for the number of SLIC segments. Our initial choice, $N=500$, was to cover an approximate area of $1300m^{2}$ for ease of interpretability and efficiency when visualised for the user (the mean average size of segments is $256\times 256\div 500=131$ — approximately 11 by 11 pixels and each pixel is $10m^{2}$ in Sentinel-2). A smaller area allows finer labelling at the cost of speed, namely projecting and labelling more points. This choice for $N$ was also optimal when considering local neighbourhood encoding using the different measures, see Table 5. For smaller $N$ values, we see that comparing segments via GLCM shows better results however, over all measures $N=500$ is the most accurate.
@@ 3.5 Graph layer ablation
3.6 Computational Complexity and Runtime Efficiency
The proposed pipeline was designed with efficiency in mind, balancing accuracy and computational cost through parameter tuning (Section 3.4). The complexity of each stage is as follows: SLIC segmentation operates in approximately $O(P)$ time, where $P$ is the number of pixels, and is highly efficient for large Sentinel-2 tiles. Fuzzy C-Means clustering introduces iterative updates but remains tractable for moderate cluster counts ($C\in\{2,8,18\}$). Using THOP, our U-Net requires 55 GFLOPs per forward pass for a 256×256 input with 12 Sentinel-2 channels. The model contains 31M trainable parameters. Graph construction scales with $O(N+K\cdot N)$, where $N$ is the number of super-pixels and $K$ the number of edges per node (here $K=8$). Message passing in the GNN is linear in the number of edges, $O(|E|)$, and was implemented with efficient batching. The GNN contains 17K parameters and requires 415 MFLOPs per forward pass. Hungarian matching introduces the highest theoretical complexity at $O(n^{3})$ for $n$ nodes per graph (approximately $n\approx 500$), but this step is applied only during similarity computation and is GPU-accelerated. Dimensionality reduction via UMAP scales well for interactive use and was precomputed for large batches.
On our test hardware (NVIDIA RTX 3090 GPU, 24 GB RAM), training the U-Net required approximately 10 hours for 600 epochs, and the GNN training completed in 1.5 hours for 300 epochs. Inference for a single tile ($42\times 42$ chips comprised of $256\times 256$ pixels) through the pipeline (segmentation, feature extraction, graph generation, graph inference, Hungarian matching, and embedding) requires approximately 23 minutes. This time is dominated by the $O(n^{3})$ Hungarian matching process (194,481 pairs of chips) which takes 20 minutes (Table 7), however is only computed once for a dataset. The application interaction achieves a frame rate update of 60 frames per second, although this is variably dependent on the amount of points selected by the user via the brushing interaction.
Table 7: Example runtime breakdown of the full processing pipeline on $1764$ chips (each $256\times 256$ pixels).
Step Time
Segmentation 2 min 20 s
Feature extraction (U-Net) 6 s
Graph generation (parallelised) 8 s
GNN 6 s
Hungarian matching, (194,481 pairs, parallelised) 20 min
Embedding (UMAP) 15 s
@@ Acknowledgements
This research was supported by EPSRC grant number EP/S021892/1 and UKHO. For the purpose of open access the authors have applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript version arising from this submission.
@@
Data Statement
Data supporting this study are included within the article and its supporting materials. The work also utilised two publicly available datasets: the Eurosat dataset by Helber et al. (Helber et al., 2019) and the Sentinel-2 Water Edges Dataset (SWED) (Seale et al., 2022). The Eurosat data is available publicly here (https://github.com/phelber/eurosat). The SWED dataset is available from https://openmldata.ukho.gov.uk/